Kyle is driving his car at 15 m/s when he noticed that he is approaching a stalled truck that is approximately 35m away . He steps on the brakes and the car stops 1.0 m away from the stalled truck . At what rate did the car slow down ? Assume that Kyle's reaction time is 2.0s . Reaction time is the time for a person to notice , think and act in response to a given situation .

1
Expert's answer
2021-11-07T19:30:17-0500

Explanations & Calculations


  • The car travels a total distance of 34m before it comes to a stop.
  • The car travels "\small d = 15ms^{-1}\times2.0s=30m" before the brakes are applied.
  • The distance it travels with brakes applied is "\small s = 34 -30 =4 m".
  • u is available, s is available and final speed v is available at the car stops.
  • Then using "\small v^2=u^2+2as", you can calculate the rate.

"\qquad\qquad
\begin{aligned}
\small 0^2&=\small 15^2+2\times a\times 4\\
\small a&=\small \bold{28.1\,ms^{-2}}
\end{aligned}"